# Adversarial Sample-Based Approach for Tighter Privacy Auditing in Final Model-Only Scenarios

### v1

@@ Abstract
Auditing Differentially Private Stochastic Gradient Descent (DP-SGD) in the final model setting is challenging and often results in empirical lower bounds that are significantly looser than theoretical privacy guarantees. We introduce a novel auditing method that achieves tighter empirical lower bounds without additional assumptions by crafting worst-case adversarial samples through loss-based input-space auditing. Our approach surpasses traditional canary-based heuristics and is effective in both white-box and black-box scenarios. Specifically, with a theoretical privacy budget of $\varepsilon=10.0$, our method achieves empirical lower bounds of $6.68$ in white-box settings and $4.51$ in black-box settings, compared to the baseline of $4.11$ for MNIST. Moreover, we demonstrate that significant privacy auditing results can be achieved using in-distribution (ID) samples as canaries, obtaining an empirical lower bound of $4.33$ where traditional methods produce near-zero leakage detection. Our work offers a practical framework for reliable and accurate privacy auditing in differentially private machine learning.
@@ 1 Introduction
Our auditing method addresses practical challenges in two scenarios within the final model setting. First, for the white-box auditing scenario, adversaries have access to the final model’s weights, as in open-source cases. We craft the worst-case adversarial sample directly using these weights. For example, with a theoretical privacy budget of $10.0$, our approach achieves an empirical lower bound of $6.68$, significantly surpassing the $4.11$ obtained with the canary approach. Secondly, for the black-box auditing scenario, only the model’s outputs are accessible, such as through APIs. We employ surrogate models to approximate the final models and use their weights to craft the worst-case adversarial sample. While adversaries’ capabilities are more limited in the black-box setting compared to the white-box setting, our approach increases the empirical lower bound from $4.11$, achieved using the canary approach, to $4.51$ for a theoretical upper bound of $10.0$. This demonstrates that our approach provides generalizable and effective privacy auditing results, even in scenarios with restricted adversarial access.
Additionally, we show that significant privacy auditing results can be achieved by using in-distribution (ID) samples from the training dataset as canaries, eliminating the need for out-of-distribution (OOD) samples. When an ID sample is used as the canary, the canary approach yields empirical privacy leakage close to zero, indicating little privacy leakage detection. In contrast, we achieve an empirical lower bound of $4.33$ even with ID canaries.
@@ 2 Preliminary
Differential Privacy (DP) (Dwork et al., 2006) is a widely used standard for preserving individual privacy in data analysis. A randomized mechanism $\mathcal{M}$ is considered $(\epsilon,\delta)$-differentially private if, for any two neighboring datasets $D$ and $D^{\prime}$ that differ by only one data point known as the canary ($c$), and for any possible output subset $O\subseteq\textit{Range}(\mathcal{M})$, the following inequality holds:
$\mathrm{Pr}[\mathcal{M}(D)\in O]\leq e^{\epsilon}\,\mathrm{Pr}[\mathcal{M}(D^{\prime})\in O]+\delta$
@@
Although DP-SGD provides a theoretical upper bound on privacy budgets ($\epsilon$, $\delta$), relying solely on privacy accounting poses challenges. First, Theoretical analyses, while crucial, are demonstrated to be conservative  (Bassily et al., 2014; Kairouz et al., 2015; Abadi et al., 2016; Mironov, 2017; Koskela et al., 2020; Gopi et al., 2021; Doroshenko et al., 2022), which can lead to overestimating the required noise and subsequently reducing the utility of the model (Nasr et al., 2021). Second, the complexity of DP-SGD training can lead to implementation errors that compromise privacy guarantees (Ding et al., 2018; Tramer et al., 2022). Privacy auditing addresses these issues by establishing empirical lower bounds on privacy, providing a more realistic assessment of privacy loss.
@@ 4 Crafting Adversarial Samples for Tighter Privacy Auditing
(a) $a_{c}$
(b) $a_{ude}$
(c) $a_{ade}$
Figure 2: The loss distributions of three adversarial samples $(a_{c},a_{ude},a_{ade})$ at a fixed privacy budget $\varepsilon=10.0$. The green distribution represents the loss outputs of models $M$, while the blue distribution represents the loss outputs of models $M^{\prime}$.
@@
Our initial loss function, $\mathcal{L}_{ude}$ (Uniform Distance Expansion), is based on the observation that the canary’s loss is naturally higher in models trained on dataset $D$ (denoted as $M_{i}$) than in models trained on $D^{\prime}$ (denoted as $M^{\prime}_{i}$). Therefore, we train the adversarial sample to ensure that models $M_{i}$ produce a higher loss and models $M^{\prime}_{i}$ produce a lower loss when evaluating $a_{w}$, aiming to maximize the distinguishability between their output distributions. The loss function $\mathcal{L}_{ude}$ is defined as:
$\mathcal{L}_{{ude}}(x_{a})=\frac{1}{N}\sum_{i=1}^{N}\left(\ell\left(M^{\prime}_{i}(x_{a}),y_{a}\right)-\ell\left(M_{i}(x_{a}),y_{a}\right)\right),$
and the corresponding minimizer, $a_{ude}$. However, as illustrated in Figure 2(b), using $\mathcal{L}_{ude}$ can lead to unnecessary training on the non-overlapping regions of the two distributions. This results in increased dispersion rather than improved separation between the distributions.
To address this issue, we propose an adaptive loss function $\mathcal{L}_{ade}$ (Adaptive Distance Expansion) that focuses solely on the overlapping regions of the loss distributions, avoiding training on already distinguishable values. This targeted approach improves the distinction between the distributions, as depicted in Figure 2(c). The adaptive loss function $\mathcal{L}_{ade}$ is defined as:
$\mathcal{L}_{{ade}}(x_{a})=\frac{1}{N}\sum_{i=1}^{N}\mathrm{ReLU}\left(\ell\left(M^{\prime}_{i}(x_{a}),y_{a}\right)-\frac{1}{N}\sum_{j=1}^{N}\ell\left(M_{j}(x_{a}),y_{a}\right)+\alpha\right),$
and the corresponding minimizer, $a_{ade}$.
In this equation, $\mathcal{L}_{ade}$ adjusts the loss based on how much each $M^{\prime}_{i}$ loss deviates from the mean loss of all models $M_{j}$. The function $\mathrm{ReLU}$ ensures that only positive deviations contribute to the loss, effectively limiting unnecessary updates when $M^{\prime}_{i}$ loss is already sufficiently separated from the mean. This allows the training process to focus on further separating less distinct distributions. Here, $\alpha$ is a tunable margin that controls the sensitivity of the separation adjustment. Increasing $\alpha$ seeks to further separate the distributions but may risk isolating only a few samples. In this work, we set $\alpha=0.2$ for all experiments. Further analysis of the impact of $\alpha$ is provided in Section 5.4.
@@ 5 Experiments
(a) MNIST
(b) CIFAR-10
Figure 3: White-box privacy auditing results on MNIST and CIFAR-10 datasets, comparing three adversarial samples: $a_{c}$, $a_{ude}$, and $a_{ade}$, across privacy budgets $\varepsilon\in\{1.0,2.0,4.0,10.0\}$.
@@ 5.1 Experiments Setup
To evaluate our auditing procedure, we conduct experiments on MNIST (LeCun et al., 1998) and CIFAR-10 (Krizhevsky, 2009) datasets with ConvNet models. We use ConvNet following the work of Annamalai and De Cristofaro (2024). We set the learning rate to $\eta=4$ for MNIST and $\eta=2$ for CIFAR-10. The training process runs for $T=100$ iterations on MNIST and $T=200$ iterations on CIFAR-10.
For models trained with DP-SGD, the accuracy on MNIST is nearly 95% for all $\varepsilon$, while the accuracy on CIFAR-10 is around 52% at $\varepsilon=10.0$. This is due to the fact that models on CIFAR-10 converge much more slowly (Annamalai and De Cristofaro, 2024); however, increasing the number of training epochs can raise the accuracy to around 70%, making it comparable to the state-of-the-art (De et al., 2022).
@@
To ensure robustness, we perform 10 independent runs for each experiment and report the average value of $\varepsilon_{\text{emp}}$ along with its standard deviation, and the theoretical upper bound for $\varepsilon$ is computed using the privacy accountant provided by Opacus (Yousefpour et al., 2021). In each run, we generate $2562N=$ 256 models–128 with canary ($M^{\prime}$) and 128 without canary ($M$). We compute empirical lower bounds using the $\mu$-GDP approach, following previous works (Nasr et al., 2023; Cebere et al., 2024), and report the lower bounds with 95% confidence intervals.
@@
5.2 White-Box Auditing in Final Model
In Figure 3, we compare the empirical privacy bounds $\varepsilon_{\text{emp}}$ between the canary sample ($a_{c}$) and our crafted worst-case samples ($a_{w}$). For $a_{c}$, we use the default canary that differentiates $D$ and $D^{\prime}$. For $a_{w}$, we employ samples generated using the loss functions $\mathcal{L}_{ude}$ and $\mathcal{L}_{ade}$, resulting in $a_{ude}$ and $a_{ade}$, respectively. Utilizing $a_{w}$ for privacy auditing consistently yields tighter empirical lower bounds across various theoretical upper bounds compared to the baseline $a_{c}$. For instance, with a theoretical privacy budget of $\varepsilon=10.0$, a canary sample $a_{c}$ produces empirical lower bounds of $4.11$ for MNIST and $0.30$ for CIFAR-10. In contrast, $a_{ude}$ and $a_{ade}$ yield estimates of $5.61$ and $6.68$ for MNIST, and $1.05$ and $4.10$ for CIFAR-10, respectively. Notably, when using $a_{c}$ on CIFAR-10, we observe almost no detectable privacy leakage. However, employing $a_{ade}$ reveals significant privacy leakage. This demonstrates $a_{ade}$’s superior effectiveness in providing tighter empirical lower bounds on privacy.
5.3 Black-box Auditing in Final Model
We also conduct privacy auditing in the final model setting when the Distinguisher has access only to the model output. In this setting, since the Distinguisher cannot directly access the weights of the final model, it is not feasible to directly train adversarial sample through final model. Instead, leveraging the fact that the Distinguisher knows both neighboring datasets $D$ and $D^{\prime}$, we apply distillation techniques to train 128 surrogate models using the output of $D^{\prime}$ for each model $M$ and $M^{\prime}$, resulting in a total of $2562N=$ 256 models. These surrogate models, distilled from the final model output in $D^{\prime}$, serve as proxies for $M$ and $M^{\prime}$ ($M_{sur}$ and $M^{\prime}_{sur}$). We then train the adversarial sample using $M_{sur}$ and $M^{\prime}_{sur}$ going through the same process as white-box auditing.
Figure 4: Black-box privacy auditing results on the MNIST dataset, comparing three auditing samples: the canary sample, $a_{ude}$, and $a_{ade}$, across privacy budgets $\varepsilon\in\{1.0,2.0,4.0,10.0\}$.
As shown in  Figure 4, although black-box auditing provides less tighter results than white-box auditing due to the attacker’s limited access, the $a_{w}$ trained by surrogate models consistently yields tighter lower bounds compared to the baseline $a_{c}=c$. Specifically, when the theoretical upper bound is $10.0$, the empirical lower bound using the baseline $a_{c}$ was $4.11$, but it increased to $4.40$ with $a_{ude}$ and $4.51$ with $a_{ade}$. Furthermore, while $a_{ade}$ provided significantly tighter auditing results than $a_{ude}$ in all privacy budgets in white-box auditing, both methods show similar lower bounds in black-box auditing. Since black-box auditing is typically limited to training interventions, crafting adversarial samples offers a promising approach for differential privacy auditing.
5.4 Additional Results
Figure 5: White-box privacy auditing results on the MNIST dataset when ordinary data is used as the canary sample, evaluated at privacy budgets $\varepsilon\in\{1.0,2.0,4.0,10.0\}$. The adversarial sample successfully audits the model, while the default canary fails.
Canary type.
Typically, the canary sample used exclusively for training $M^{\prime}$ is an out-of-distribution (OOD) sample that differs significantly from the training dataset $D$. Without an OOD canary, obtaining tighter lower bounds becomes challenging, as the diffrence between $M$ and $M^{\prime}$ is minimal. However, leveraging a worst-case adversarial sample enables significantly tighter lower bounds, even when using in-distribution (ID) canary.
As shown in Figure 5, audits using a standard ID canary sample yield zero empirical privacy leakage estimates ($\varepsilon_{\text{emp}}$) across all privacy budgets when applying the baseline attack $a_{c}$. In contrast, adversarial samples improve auditing performance significantly. At $\varepsilon=10.0$, the empirical lower bounds rise to 4.05 for $a_{ude}$ and 4.33 for $a_{ade}$, demonstrating the effectiveness of $a_{w}$ in exposing privacy leakage. Furthermore, $a_{ade}$ consistently achieves slightly tighter bounds than $a_{ude}$ across all privacy budgets, highlighting the robustness of adaptive attacks in privacy auditing. These results emphasize the importance of leveraging adversarial sample when auditing with ID canary.
Figure 6: Privacy auditing results on the MNIST dataset using both in-distribution (ID) and out-of-distribution (OOD) data as the canary samples (blank, mislabel, clip-bkd), evaluated at a privacy budget of $\varepsilon=10.0$.
Additionally, adversarial samples are effective across diverse canary types, including ID samples, blank samples, mislabeled samples and a clip-bkd samples. The mislabeled image is an OOD sample intentionally assigned the wrong label, and the clip-bkd image, crafted as proposed by  Jagielski et al. (2020), target the direction of least variance to create a robust attack against the gradient clipping norm.
As shown in Figure 6, $a_{ade}$ consistently achieves tighter lower bounds compared to the baseline ($a_{c}$) across all canary types. Specifically, for ID canary, blank image, mislabeled image, and clip-bkd image, our method achieves $\varepsilon_{\text{emp}}=3.03$, $3.94$, $3.96$, and $3.91$, respectively, significantly outperforming the baseline values of $\varepsilon_{\text{emp}}=0.00$, $1.29$, $0.93$, and $1.56$. These results highlight the robustness of our approach in privacy auditing, demonstrating its effectiveness across ID and OOD canary.
(a) MNIST
(b) CIFAR-10
Figure 7: Privacy auditing results on the MNIST and CIFAR-10 datasets using worst-case initialization with adversarial sample, evaluated across privacy budgets $\varepsilon\in\{1.0,2.0,4.0,10.0\}$.
Worst case initialization.
Recent work (Annamalai and De Cristofaro, 2024) leverages the insight that minimizing the gradients of non-canary samples enables tight auditing in DP-SGD, aiming to achieve this in black-box auditing of the final model by pre-training on an auxiliary dataset matching the training data distribution. These crafted worst-case initial model parameters are then used as the model’s initial weights, replacing random initialization to obtain tighter lower bounds in a black-box setting in final model.
In this section, we present our results on tighter privacy auditing for the final model trained on MNIST and CIFAR-10 datasets by combining the worst-case model parameters ($\theta_{\text{worst}}$) with a worst-case adversarial sample ($a_{w}$). As illustrated in Figure 7, using the canonical attack $a_{c}$ with the average model parameters $\theta_{\text{avg}}$ yielded lower bounds of $4.11$ on MNIST and $0.41$ on CIFAR-10. When we switched to the worst-case model parameters $\theta_{\text{worst}}$ while using the same attack $a_{c}$, these lower bounds increased to $6.83$ and $5.31$, respectively, indicating tighter estimates of privacy leakage. Further improvements were achieved by employing the adaptive attack $a_{\text{ade}}$ with $\theta_{\text{worst}}$, which resulted in even higher lower bounds of $9.9$ on MNIST and $9.93$ on CIFAR-10. Notably, across all theoretical privacy budgets, the auditing results using $\theta_{\text{worst}}$ with $a_{\text{ade}}$ closely aligned with the theoretical upper bounds. These findings demonstrate that we achieved tight privacy auditing, reinforcing the argument that privacy amplification in the final model setting does not occur with general non-convex loss.
Figure 8: Auditing models trained with varying gradient clipping norms $C$, evaluated at privacy budgets $\varepsilon=10.0$.
Clipping norm sensitivity.
Based on the experimental results, we analyze the impact of varying gradient clipping norms on the empirical privacy leakage estimates ($\varepsilon_{\text{emp}}$). Figure 8 presents the outcomes for three attack scenarios: $a_{c}$, $a_{ude}$, and $a_{ade}$.
When the gradient clipping norm is small ($C=0.1$), the privacy leakage estimates are 5.90, 7.81 and 7.85 for $a_{c}$, $a_{ude}$ and $a_{ade}$, respectively, indicating tight audits due to a high signal-to-noise ratio. As the clipping norm increases to $C=1.0$, $\varepsilon_{\text{emp}}$ decreases across all scenarios; however, the rate of decrease varies. Notably, $a_{ade}$ is less affected compared to $a_{c}$ and $a_{ude}$, retaining a relatively high estimate of 4.24. This suggests that $a_{ade}$ is more robust and capable of effectively auditing privacy leakage even with larger clipping norms. At a higher clipping norm ($C=10.0$), $\varepsilon_{\text{emp}}$ decreases significantly for all attacks, with values of 0.00, 0.59, and 0.87 for $a_{c}$, $a_{ude}$, and $a_{ade}$, respectively. This reduction is attributed to excessive noise overwhelming the useful signal, resulting in looser audits.
@@ 5.4 Additional Results
Table 1: Privacy auditing results on the CIFAR-10 dataset for various values of $\alpha$ with a theoretical privacy budget of $\varepsilon=4.0$.
$\alpha$ 0.05 0.1 0.15 0.2 0.25
$\epsilon_{emp}$ $1.90_{\pm 0.51}$ $1.84_{\pm 0.32}$ $1.83_{\pm 0.34}$ $1.87_{\pm 0.21}$ $2.03_{\pm 0.24}$
These results highlight the trade-offs between gradient clipping norms and the tightness of privacy audits, emphasizing that adaptive attacks like $a_{ade}$ can provide more effective auditing even under higher clipping norms.
$\alpha$ sensitivity.
Although the worst-case adversarial sample demonstrates tighter privacy auditing performance, it relies on selecting an appropriate value for the hyperparameter $\alpha$. To assess the effect of $\alpha$, we trained adversarial samples using the $\mathcal{L}_{ade}$ with $\alpha$ set to various values: $0.05$, $0.1$, $0.15$, $0.2$, and $0.25$. As shown in Table 1, the $\mathcal{L}_{ade}$ exhibits similar auditing performance across these values, indicating that our loss function is robust to the choice of $\alpha$, which is advantageous. Empirically, we find that $\alpha$ values between $0.02$ and $0.5$ do not result in significant performance degradation.

### v2

@@ Abstract
Auditing Differentially Private Stochastic Gradient Descent (DP-SGD) in the final model setting is challenging and often results in empirical lower bounds that are significantly looser than theoretical privacy guarantees. We introduce a novel auditing method that achieves tighter empirical lower bounds without additional assumptions by crafting worst-case adversarial samples through loss-based input-space auditing. Our approach surpasses traditional canary-based heuristics and is effective in final model-only scenarios. Specifically, with a theoretical privacy budget of $\varepsilon=10.0$, our method achieves empirical lower bounds of $4.914$, compared to the baseline of $4.385$ for MNIST. Our work offers a practical framework for reliable and accurate privacy auditing in differentially private machine learning.
@@ 1 Introduction
Our auditing method addresses practical challenges in the final model setting. Adversaries have access to the final model’s weights, as in open-source cases. We craft the worst-case adversarial sample directly using these weights. For example, with a theoretical privacy budget of $10.0$, our approach achieves an empirical lower bound of $4.914$, significantly surpassing the $4.385$ obtained with the canary approach.
@@ 2 Preliminary
Differential Privacy (DP) (Dwork et al., 2006) is a widely used standard for preserving individual privacy in data analysis. A randomized mechanism $\mathcal{M}$ is considered $(\varepsilon,\delta)$-differentially private if, for any two neighboring datasets $D$ and $D^{\prime}$ that differ by only one data point known as the canary ($c$), and for any possible output subset $O\subseteq\textit{Range}(\mathcal{M})$, the following inequality holds:
$\mathrm{Pr}[\mathcal{M}(D)\in O]\leq e^{\varepsilon}\,\mathrm{Pr}[\mathcal{M}(D^{\prime})\in O]+\delta$
@@
Although DP-SGD provides a theoretical upper bound on privacy budgets ($\varepsilon$, $\delta$), relying solely on privacy accounting poses challenges. First, Theoretical analyses, while crucial, are demonstrated to be conservative  (Bassily et al., 2014; Kairouz et al., 2015; Abadi et al., 2016; Mironov, 2017; Koskela et al., 2020; Gopi et al., 2021; Doroshenko et al., 2022), which can lead to overestimating the required noise and subsequently reducing the utility of the model (Nasr et al., 2021). Second, the complexity of DP-SGD training can lead to implementation errors that compromise privacy guarantees (Ding et al., 2018; Tramer et al., 2022). Privacy auditing addresses these issues by establishing empirical lower bounds on privacy, providing a more realistic assessment of privacy loss.
@@ 3 DP-SGD Auditing procedure
(a) $a_{c}$
(b) $a_{L_{2}}$
(c) $a_{F}$
(d) $a_{BD}$
Figure 2: The loss distributions of the canary sample $a_{c}$ and three adversarial samples $(a_{L_{2}},a_{F},a_{BD})$ in the training set at a fixed privacy budget $\varepsilon=1.0$. The green distribution represents the loss outputs of models $M$, while the blue distribution represents the loss outputs of models $M^{\prime}$.
@@ 4 Crafting Adversarial Samples for Tighter Privacy Auditing
Distance-Based Separation.
@@
Intuitively, the canary’s loss tends to be higher for models trained on the dataset $D$ (denoted by $M_{i}$) than for those trained on $D^{\prime}$ (denoted by $M^{\prime}_{i}$). To exploit this discrepancy, we train an adversarial sample $a_{w}$ that encourages $M_{i}$ to produce a higher loss and $M^{\prime}_{i}$ to produce a lower loss, thereby maximizing the distinguishability between their output distributions. As a straightforward approach, we utilize the L2 loss, $\mathcal{L}_{L_{2}}$, to widen this gap:
$\mathcal{L}_{L_{2}}(x_{a})=\frac{1}{N}\sum_{i=1}^{N}\Bigl{[}\ell\bigl{(}M^{\prime}_{i}(x_{a}),y_{a}\bigr{)}\;-\;\ell\bigl{(}M_{i}(x_{a}),y_{a}\bigr{)}\Bigr{]},$
and denote the adversarial sample generated by this objective as $a_{L_{2}}$.
Distributional Separation.
Additionally, we apply two distribution-based losses to maximize the separation between $M^{\prime}$ and $M$. The first is the Bhattacharyya loss, $\mathcal{L}_{BD}$, which is derived from the Bhattacharyya distance, a measure of overlap between two probability distributions. When the outputs of $M_{i}(x_{a})$ and $M^{\prime}_{i}(x_{a})$ can be approximated as univariate Gaussians $\mathcal{N}(\mu_{M_{i}},\sigma_{M_{i}}^{2})$ and $\mathcal{N}(\mu_{M^{\prime}_{i}},\sigma_{M^{\prime}_{i}}^{2})$, the Bhattacharyya distance between them is:
$D_{B}\bigl{(}\mathcal{N}(\mu_{M_{i}},\sigma_{M_{i}}^{2}),\mathcal{N}(\mu_{M^{\prime}_{i}},\sigma_{M^{\prime}_{i}}^{2})\bigr{)}=\frac{(\mu_{M_{i}}-\mu_{M^{\prime}_{i}})^{2}}{4\bigl{(}\sigma_{M_{i}}^{2}+\sigma_{M^{\prime}_{i}}^{2}\bigr{)}}\;+\;\frac{1}{2}\ln\!\Bigl{(}\frac{\sigma_{M_{i}}^{2}+\sigma_{M^{\prime}_{i}}^{2}}{2\,\sigma_{M_{i}}\,\sigma_{M^{\prime}_{i}}}\Bigr{)}.$
To minimize this overlap, we define $\mathcal{L}_{BD}$ as the negative of the Bhattacharyya distance:
$\mathcal{L}_{BD}=-\frac{1}{N}\sum_{i=1}^{N}D_{B}\bigl{(}\mathcal{N}(\mu_{M_{i}},\sigma_{M_{i}}^{2}),\mathcal{N}(\mu_{M^{\prime}_{i}},\sigma_{M^{\prime}_{i}}^{2})\bigr{)}.$
Minimizing $\mathcal{L}_{BD}$ thereby pushes the two distributions apart, enhancing the distinguishability between $M$ and $M^{\prime}$.
Additionally we adopt Fisher’s Linear Discriminant Analysis (LDA). LDA maximizes the ratio of the squared difference between class means to the sum of their variances, thereby encouraging a clear separation between distributions. Following this principle, we define the Fisher Discriminant Loss as:
$\mathcal{L}_{F}=-\frac{1}{N}\sum_{i=1}^{N}\frac{(\mu_{M^{\prime}_{i}}-\mu_{M_{i}})^{2}}{\sigma_{M^{\prime}_{i}}^{2}+\sigma_{M_{i}}^{2}}.$
By normalizing the squared difference of the means by the sum of their variances, $\mathcal{L}_{F}$ remains stable and effectively preserves the separation between $M$ and $M^{\prime}$. We denote the adversarial samples produced by minimizing the Bhattacharyya loss and the Fisher Discriminant Loss as $a_{BD}$ and $a_{F}$, respectively; both aim to maximize the distributional gap between $M$ and $M^{\prime}$.
Data Splitting for Adversarial Samples.
To mitigate the risk of information leakage when generating adversarial samples, we adopt a strict data-splitting protocol. Specifically, we derive the adversarial samples exclusively from the training models and then measure the empirical lower bounds of differential privacy on an independent validation models. This separation ensures that any potential leakage from the training process does not influence the subsequent evaluation, thereby yielding a more reliable assessment of the privacy guarantees. Overall, this strategy effectively decouples the adversarial sample generation from the empirical DP evaluation, reducing bias and strengthening the credibility of our results.
@@ 5.1 Experiments Setup
To evaluate our auditing procedure, we conduct experiments on the MNIST (LeCun et al., 1998) dataset using ConvNet models. Following the architecture used in Annamalai and De Cristofaro (2024), we train the models with a learning rate of $\eta=4$ for $T=100$ iterations. For models trained with DP-SGD, the accuracy on MNIST remains consistently close to 95% across all values of $\varepsilon$, indicating the stability of the training process under differential privacy constraints.
@@
To ensure robustness, we perform 10 independent runs for each experiment and report the average value of $\varepsilon_{\text{emp}}$ along with its standard deviation, and the theoretical upper bound for $\varepsilon$ is computed using the privacy accountant provided by Opacus (Yousefpour et al., 2021). In each run, we generate $10242N=$ 1024 models–512 with canary ($M^{\prime}$) and 512 without canary ($M$). We compute empirical lower bounds using the $\mu$-GDP approach, following previous works (Nasr et al., 2023; Cebere et al., 2024), and report the lower bounds with 95% confidence intervals.
@@
Table 1: Privacy auditing results on the MNIST dataset with various adversarial samples (Canary, L2, Fisher, Bhattacharyya) in theoretical privacy budget of $\varepsilon=\{1.0,2.0,4.0,10.0\}$.
Loss $\varepsilon=1.0$ $\varepsilon=2.0$ $\varepsilon=4.0$ $\varepsilon=10.0$
Canary $0.212_{\pm 0.177}$ $0.594_{\pm 0.107}$ $1.696_{\pm 0.207}$ $4.385_{\pm 0.284}$
L2-Distance $0.215_{\pm 0.119}$ $0.646_{\pm 0.142}$ $1.722_{\pm 0.248}$ $4.788_{\pm 0.255}$
Fisher $\underline{0.226}_{\pm 0.112}$ $\underline{0.648}_{\pm 0.116}$ $\textbf{1.823}_{\pm 0.335}$ $\textbf{4.914}_{\pm 0.268}$
Bhattacharyya $\textbf{0.232}_{\pm 0.141}$ $\textbf{0.661}_{\pm 0.120}$ $\underline{1.785}_{\pm 0.297}$ $\underline{4.854}_{\pm 0.268}$
5.2 Auditing Results in Final Model
To evaluate the effectiveness of adversarial sample selection in privacy auditing, we compare the empirical privacy lower bounds $\varepsilon_{\text{emp}}$ obtained using the default canary sample ($a_{c}$) and our crafted worst-case adversarial samples ($a_{w}$) in Table 1. Specifically, $a_{c}$ refers to the default canary that differentiates $D$ and $D^{\prime}$, whereas $a_{w}$ consists of adversarial samples generated using the loss functions $\mathcal{L}_{L_{2}}$, $\mathcal{L}_{BD}$, and $\mathcal{L}_{F}$.
Figure 3: Privacy auditing results on the MNIST dataset using $\{128,256,384,512\}$ models.
Across different theoretical privacy budgets $\varepsilon$, employing $a_{w}$ consistently yields tighter empirical lower bounds than the baseline $a_{c}$. As shown in Table 1, for $\varepsilon=1.0$, the default canary achieves an empirical bound of 0.212, while Fisher and Bhattacharyya loss-based samples improve it to 0.226 and 0.232, respectively. This trend persists at higher privacy budgets, where Fisher-based samples achieve the highest empirical bound of 4.914 at $\varepsilon=10.0$, outperforming the canary baseline (4.385). Notably, both Fisher and Bhattacharyya-based adversarial samples consistently provide the tightest privacy lower bounds across all settings, confirming their effectiveness in maximizing privacy distinguishability.
5.3 Effect of Number of Models
As shown in Figure 3, adversarial samples consistently achieve higher auditing performance as the number of models grows, rising from 4.44 with 128 models to 4.91 with 512 models for $a_{F}$. This improvement arises because, as the number of models grows, the distributions of training-model and evaluation-model losses become more similar. This closer alignment enables adversarial samples to be more effectively optimized, ultimately leading to enhanced auditing outcomes.